OpenStudy (anonymous):

Find the distance between the points given. Simplify all irrational answers. (2, 3) and (-10, 12)

OpenStudy (saifoo.khan):

Do you know the distance formula?

OpenStudy (anonymous):

yup

OpenStudy (saifoo.khan):

Do you know how to use it?

OpenStudy (anonymous):

yep

OpenStudy (lgbasallote):

so what's the problem?

OpenStudy (saifoo.khan):

Did you tried to solve?

OpenStudy (anonymous):

yup i did

OpenStudy (saifoo.khan):

Got the answer?

OpenStudy (anonymous):

and i got

OpenStudy (saifoo.khan):

The answer you got was.. ?

OpenStudy (anonymous):

sqrt 145

OpenStudy (anonymous):

but it wasnt the correct answer,,,

OpenStudy (saifoo.khan):

You should get sqrt225

OpenStudy (anonymous):

use the equation: \[\sqrt{(y _{1}-y _{2})^2+(x _{1}-x _{2})^2}\]

OpenStudy (anonymous):

how? @saifoo.khan

OpenStudy (lgbasallote):

(-10 - 12)^2 + (12 - 3)^2 = 144 + 81

OpenStudy (lgbasallote):

that's (-10 - 2)^2

OpenStudy (saifoo.khan):